OpenStudy (anonymous):

What are the sine, cosine, and tangent of 2 pi over 3 radians

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

(2 pie)/ 3 = 120 degree here pie = 180

OpenStudy (anonymous):

so tan is also 0?

OpenStudy (anonymous):

sin (120) = sin(180-60) = sin60 = sqrt(3)/2

OpenStudy (anonymous):

and proceed same way to cosine and tangent

OpenStudy (anonymous):

wow i was thinking 3pi/2 all this time wtffffff

OpenStudy (anonymous):

my mistake

OpenStudy (anonymous):

would cosine be + or - 1/2?

OpenStudy (anonymous):

@Yahoo! @magepker728

OpenStudy (anonymous):

cosine in 2 q2 is negative

OpenStudy (anonymous):

thank you!